Please. I need help evaluating this problem.
p+q/5 for p=44 and q=1 (simplify the answer)
Thank you

Answer by stanbon(75887)

p+q/5 for p=44 and q=1
If you mean p+(q/5) the answer is 44+(1/5) = 44.2
-------------------
If you mean (p+q)/5 the answer is (44+1)/5 = 45/5 = 9
--------------------
Notice how important the parentheses are.
That is why we recommend you use them when posting problems.
---------------------
cheers,
Stan H